can you please answer this??
(Explain the following about problem 9.31.
a. Why do you use the electron affinity of F twice, but two separate ionization energies for Mg?
b. Why is the lattice energy of this so much larger than the 788kJ/mol that you calculated for NaCl)


KNOWING THAT PROBLEM 9.31 is :
Use the following information to calculate the lattice energy of MgF2. This is actually Problem 9.31
Mg(s)  Mg(g) ΔH = 148 kJ
F2(g)  2 F(g) ΔH = 159 kJ
Mg(g)  Mg+(g) + e- ΔH = 738 kJ
Mg(g)  Mg2+ (g) + e- ΔH = 1450 kJ
F(g) + e-  F-(g) ΔH = -328 kJ
Mg(s) + F2(g)  MgF2(s) ΔH = -1123kJ
Answer #1

Solution : The overall reaction is a Mg(8) + F78) MgGz (8) DH=-1123 k3 /mol. @ since we are dealing with 2f atoms (f2), and tThus, we have AH47 148 + 9+138 +140-32880 = -1123 kJ/mole = 1839 tu or o=11839 + 1123) kJ / mole =F-2962 kJ/moll This is muchnote- in born lande equation the symbols have their usual meanings, and n= born exponent.